[Ncep.nhc.nco_contacts] TSB Morning Rounds - Fri Aug 18, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Fri Aug 18 14:19:56 UTC 2017


Here is today's summary of NHC computer operations:
--- Monica

----------

TAFB:
AC Unit 3 in server room alarming showing failure temp 80F. TSB will  
investigate.
HSU:
1) The ridges workstation would not allow Lixion to do the graphics on the  
intermediate advisory. It kept crashing (exiting NMAP2 with a bad request  
error). He ended up doing the product from the backup workstation. This is  
probably related to the graphics card issues (see below), but TSB will keep  
a watch on it.
2) The phone system was upgraded yesterday and the NCO ticket has been  
closed.
Other items:
1) GOES-16 will be placed in continuous full disk mode for testing August  
19th (at 21Z) until the 21st (at 2Z), 2017 for approximately 29 hours. A  
full disk image will be distributed every 5 minutes in this mode, and both  
the CONUS and Meso sectors will not be available.
2) A GOES-16 solar eclipse monitoring scan will be performed on August  
21st, from 1645-2000 UTC resulting in an outage for MESO-1. MESO-2 will  
still be available during this time.
3) Joe Zajic will be updating the GOES-16 processing code Monday (8/21)  
morning. The switch should be transparent, but there may be a minor outage  
for a few minutes.
4) Monitor upgrade work on the operations floor continues.
5) NAWIPS Workstations Video Card Upgrades: NCO has upgraded the TAFB  
Pacific satellite workstation so far, additional upgrades are in the works.

NHC RFCs:
None

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170818/8992f6ff/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list